A straight line makes on the coortinate axes positive intercepts whose sum is 7. If the line pass through the point (-3,7 ) find its equation. Plzz do this question

Grade:10

1 Answers

Arun
6 years ago
Dear Vaibhav
 
Let line
x/a + y/b = 1
Now
a + b = 7
Also
-3/a + 7/b = 1
-3b + 7a = ab
Now 
Solve both equation and find a and b
Hence you will get your equation of line.
 
Regards
Arun (askIITians forum expert)